OpenStudy (anonymous):

PLEASE HELP!!! MEDAL WILL BE REWARDED!!!! For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 8 + . . . + 2n = 2(n+1) - 2

OpenStudy (anonymous):

P(n) = 2(n+1)-2 thefore P(k) = 2(k+1)-2 and P(k+1) = 2(k+1-1)-2